# RLJP: Legal Judgment Prediction via First-Order Logic Rule-enhanced with Large Language Models

### v1

@@ 1 Introduction
The current methods of LJP mainly consist of two approaches: modeling legal knowledge and retrieving judicial precedents or legal knowledge. (1) Methods modeling legal knowledge typically use neural networks to model legal provisions and extract semantic associations between case facts and legal knowledgeXu et al. (2020, 2024). While these models can efficiently extract relevant legal knowledge for LJP relying on semantic similarity, they overlook the intrinsic logic of cases. (2) Other methods leveraged semantic techniques to search for related judicial precedents or legal knowledgeWu et al. (2023a); He et al. (2024). Judicial precedents and legal knowledge as references to enhance the model’s understanding of the case circumstances, but they tend to study similar cases or knowledge instead of modeling the logical reasoning process. In summary, the above two types of methods rely on textual and semantic matching but ignore capturing the logic of judging legal cases, where logic in reasoning is crucial in legal judgment.
To address the aforementioned issues, researchers applied legal judgment logic to enhance reasoning in traditional deep learning models or large language models (LLMs) for LJP. Some researchers employ deep learning models to extract logical premises from complex cases Yue et al. (2021, 2024) or handle the legal judgment using rules defined by experts Gan et al. (2021). The integration of legal judgment logic has improved the reasoning capabilities of deep learning models. However, the ability of deep learning models is still limited by their model capacity and training data scale. As LLMs have remarkable abilities in text comprehension and logical reasoning, researchers combined judgment logic with LLMs for LJP. ResearchersDeng et al. (2023) use LLMs to retrieve relevant legal provisions and extract four types of criminal elements from case facts for judgment based on the classical legal logic syllogism. Prompts constructed by keyword structure from legal experts’ determination guide LLMs to classify legal casesIzzidien et al. (2024). In summary, while existing methods using legal logic effectively analyze cases through established rules, their logic rigidity limits adaptation to case-specific contexts in complex scenarios.
@@ 2 Related Work
LJP aims to predict judicial outcomes by employing traditional models or LLMs to analyze legal cases He et al. (2023); Deng et al. (2024); Zhang et al. (2024). Broadly, LJP can be divided into two main approaches: one based on semantic similarity Liu et al. (2024a) and the other guided by judgment logic Cheng et al. (2024).
@@ 2.1 LJP Assisted by Semantic Similarity
The semantic similarity-assisted method for LJP involves modeling domain knowledge and applying retrieval techniques Zhang and Dou (2023).
The former mainly utilizes neural networks to model legal knowledge Yue et al. (2021); Yao et al. (2024) for feature extraction and to establish semantic associations among the knowledge. To improve the accuracy of crime prediction, the prevailing methods primarily focus on refining the attention mechanism He et al. (2023); Yu and Qiu (2023); He et al. (2025). Additionally, some methods infuse domain-specific legal knowledge into LLMs Wang et al. (2024); Li et al. (2025), thereby enhancing the models’ capability to identify and leverage critical information. Regarding confusion issue in criminal charges, Zhang et al. (2023) and Gan et al. (2023) utilized contrastive learning to capture fine-grained distinctions. Using retrieval techniques refers to utilizing search engines Wu et al. (2023a); Yue et al. (2023); Liu et al. (2024b) to obtain external knowledge from semantically similar judicial precedents. He et al. (2024) introduced the retrieval technology into the SimCourt multi-agent framework. Qin et al. (2024) constructed a hierarchical semantic ID for the retrieved candidate documents.
@@ 2.2 LJP Assisted by Case Judgment Logic
The legal judgment logic-enhanced method for LJP leverages models to summarize the knowledge embedded in case details Chen et al. (2020); Yang et al. (2022) and extract the underlying reasoning logic.
Existing methods focus on using the entire factual description Chai et al. (2020); Hong and Chang (2023) to generate judgment results, overlooking the actual judicial scenario where judges consider various criminal circumstances Wu et al. (2022) to determine the verdict and sentencing. Consequently, Yue et al. (2021) proposed a legal judgment framework, NeurJudge, which can divide facts and the prediction results of intermediate subtasks into different scenarios. Yue et al. (2024) further proposed NeurJudge+, which integrates the semantic labels of accusations and legal provisions into facts. Gan et al. (2021) injected legal knowledge as an additional logical rule into the neural network. Deng et al. use LLMs to retrieve relevant legal provisions and extract four types of criminal elements from case facts for judgment based on the classical legal syllogism. Prompts constructed by keyword structure from legal experts’ determination guide LLMs to classify legal casesIzzidien et al. (2024).
@@ 3.1 Task Definition
  • Judgment is the final decision made by a judge in a legal case based on the facts and the judgment rule. It typically consists of the law article, the charge, and the prison term. We represent the judgment of a case as $j=(article,charge,prison\_term)$, where $(article,charge,prison\_term)$ refers to the labels of provisions, accusation, and prison term, respectively.
@@
Problem 2 Legal Judgment Prediction. Given the case fact $f$, our task is to analyze $f$ based on $rule$, then predict judgment $j=(article,charge,prison\_term)$.
@@ 3.3 Rules Initialization
Each FOL judgment rule $Rule:A\to C$ comprises an antecedent $A$ and a consequent $C$. The content of $A$ is determined by the circumstance logic. And $C$ is constructed by one or two legal judgment labels, including $article$, ($article$, $charge$), and ($article$, $prison\_term$).
@@
(3) Construct FOL judgment rules. The purpose of this step is to standardize the format of rules to facilitate subsequent optimization and reasoning processes. The judgment rules are constructed using the antecedent $A$ and the consequent $C$. $Rule:A\to C$ in this step. The antecedent $A$ comprises multiple FOL symbols and logical connectives ($\vee$, $\wedge$, and $\neg$). The consequent $C$ is the same as LJP labels, including legal provisions $article$, (legal provisions $article$, accusation $charge$), and (legal provisions $article$, prison terms $prison\_term$).
@@ 3.4 Confusion-Aware Rule Optimize Engine
0:  $Precedents$: Precedent set;
0:  $target$: Target label which is the consequence of the rules;
0:  $RULE$: Initial rule set ;
0:  $DefinedScore$: A predefined score threshold serves as the stopping criterion for terminating the iterative updating process;
0:  $Num$: The required number of confusable cases.
1:  $S_{postive}\leftarrow Precedents[target]$
2:  $S_{other}\leftarrow Precedents[other]$
3:  $S_{negative}\leftarrow RankSimilarity(S_{postive},S_{other},Num)$
@@
5:  $r_{0}\leftarrow RULE[target]$
6:  $MaxScore\leftarrow 0,MaxPointer\leftarrow Null$
17:  $R_{tree}\leftarrow[r_{0}]$
@@
9:     if Is_Evaluate($r$then
10:        ($score$, reasoning records) $\leftarrow history(r)$
@@
12:        ($score$, reasoning records) $\leftarrow Evaluate(V_{target},r)$
@@
22:     $R^{\prime}\leftarrow CACL(R^{*},$ reasoning records ))))
@@ 3.4.1 Construction of the Confusable Cases Validation Set
To find confusable cases for legal judgment $target$, we first find the case set, whose legal judgment label is $target$, denoted as $S_{positive}$, and the case set for other legal judgment labels, denoted as $S_{other}$. Then, to asses the similarity score between $S_{positive}$ and $S_{other}$, we convert each case fact set into fixed-dimensional embedding vectors using the BGE model: $E_{t}=\left\{E_{t}^{1},E_{t}^{2},...,E_{t}^{m}\right\}$ for $S_{positive}$ and $E_{o}=\left\{E_{o}^{1},E_{o}^{2},...,E_{o}^{n}\right\}$ for $S_{other}$, $E_{t}^{i},E_{o}^{i}\in R^{d}$. Then, we compute the cosine similarity between $E_{t}$ and $E_{o}$. The results are organized into a similarity matrix $A$, where $A(i,j)=\frac{E_{t}^{i}\cdot E_{o}^{i}}{\left\|E_{t}^{i}\right\|\left\|E_{o}^{i}\right\|}$ represents the similarity score of two cases $S_{positive}^{i}$ and $S_{other}^{j}$.
The higher the similarity score between cases in different judgment labels means that they contain similar crime circumstances, the more likely they are to be confusable for legal judgment. For this reason, we sort the similarity matrix $A$ in descending order based on the similarity scores of each row, and the first column is extracted to obtain a set of negative sample cases for judgment $target$: $S_{negative}\in S_{other}$. Finally, the confusable case set $V_{target}$ can be constructed: $V_{target}=S_{negative}\cup S_{positive}$.
@@ 3.4.2 Optimization Tree Splitting
0:  $R^{*}$: Current rule (Anchor);
0:  $Correct(R^{*})$: Set of positive samples (valid reasoning records);
0:  $Incorrect(R^{*})$: Set of negative samples (invalid reasoning records).
1:  $\bigtriangledown{keep}\leftarrow LLM(R^{*},Correct(R^{*}))$
2:  $\bigtriangledown{imp}\leftarrow LLM(R^{*},Incorrect(R^{*}))$
3:  $\bigtriangledown{R}\leftarrow LLM(\bigtriangledown{keep},\bigtriangledown{imp})$
14:  $R^{\prime}\leftarrow LLM(\bigtriangledown{R},R^{*})$
@@ 4.1 Settings
Datasets. Following previous LJP works (Zhong et al., 2018; Xu et al., 2020; Yue et al., 2021; Wu et al., 2023b), we conduct experiments on datasets CAIL2018  [^2] https://github.com/china-ai-law-challenge and CJO22(Wu et al., 2023b). For the CAIL2018 dataset, we randomly divide it into training, validation, and test sets with a ratio of 8:1:1Wu et al. (2023b). For the CJO22 dataset, following Wu et al. (2023b), we use it solely as an additional test set. The table in Appendix F presents the details of two datasets.
Evaluation Metrics. We use Accuracy (Acc), Macro-Precision (Ma-P), Macro-Recall (Ma-R), and Macro-F1 (Ma-F) as the evaluation metrics following previous LJP works (Zhong et al., 2018; Xu et al., 2020; Yue et al., 2021; Wu et al., 2023b).
Baselines. We compare with: (1) CNN(LeCun et al., 1989) use different kernel convolutional operations to extract text features for classification; (2) BERT(Devlin et al., 2019) can be easily fine-tuned on downstream tasks such as LJP; (3) TopJudge(Zhong et al., 2018) employs multi-task learning and captures the dependencies among the three sub-tasks in LJP; (4) R-FormerDong and Niu (2021) formalizes LJP as a node classification problem over a global consistency graph; (5) LADANXu et al. (2020) uses graph distillation to extract discriminative features of the fact; (6) NeurJudge(Yue et al., 2021) splits the fact description into different parts for making predictions; (7) EPMFeng et al. (2022) locates event-related information essential for judgment; (8) CTM(Liu et al., 2022) uses case triple modeling from contrastive case relations; (9) PLJP(Wu et al., 2023b) uses the collaboration between LLMs and domain-specific models for LJP; (10) Llama3(Wu et al., 2023b), which is Meta’s advanced open-source language model; (11) D-LADAN(Xu et al., 2024) solves LJP confusion via graph and memory mechanisms.
Implementation Details. We directly adopted Bert-base-chinese [^3] https://huggingface.co/google-bert/bert-base-chinese for candidate labels, Gpt-4o for rules optimization, and Llama3-chineseZhichen Zhang (2024) for other steps because the language of the cases in the data set is Chinese. For the length limit, we use three precedents in Section 3.3. For the parameter settings of the baseline, we strictly follow the original paper. More details are in Appendix G.
@@ 4.2 Main Results
Method Law Article Charge Prison Term
Acc Ma-P Ma-R Ma-F Acc Ma-P Ma-R Ma-F Acc Ma-P Ma-R Ma-F
CNN(LeCun et al., 1989) 80.50 40.10 38.33 38.49 87.52 88.23 88.31 88.17 34.42 32.22 30.53 31.05
BERT(Devlin et al., 2019) 82.77 36.82 35.94 35.82 89.10 90.10 89.48 89.63 40.00 37.53 33.66 33.58
Roberta(Zhong et al., 2018) 83.08 48.09 44.25 44.87 90.30 91.02 90.97 90.94 40.84 38.62 38.55 38.50
TopJudge(Zhong et al., 2018) 80.46 40.96 40.96 38.24 87.31 88.68 87.84 88.20 35.54 33.55 31.08 32.00
R-FormerDong and Niu (2021) 87.82 56.13 56.57 55.81 91.54 91.61 91.96 91.58 40.70 36.09 36.76 35.04
LADANXu et al. (2020) 82.82 42.57 39.00 40.71 88.09 90.12 88.82 89.47 38.03 33.66 30.08 31.77
Neurjudge(Yue et al., 2021) 76.91 55.95 52.92 53.56 82.13 82.71 82.30 82.36 33.53 36.46 37.26 36.53
EPMFeng et al. (2022) 85.80 49.08 45.76 47.32 91.20 90.81 89.99 90.46 40.25 37.96 37.00 37.34
CTM(Liu et al., 2022) 84.72 46.46 44.83 45.10 90.28 90.34 88.08 86.30 39.56 38.66 38.02 37.84
Llama3-chineseZhichen Zhang (2024) 2.62 4.41 1.66 1.11 22.63 22.53 9.58 10.04 12.00 14.91 19.04 10.53
PLJP(Bert)(Wu et al., 2023b) 87.07 58.81 57.29 56.63 94.99 92.12 91.10 91.33 48.72 42.64 36.80 35.43
D-LADAN(Xu et al., 2024) 91.05 56.37 54.91 58.38 90.82 66.26 62.51 62.08 38.31 27.48 26.01 25.00
RLJP(ours) 91.27 85.72 91.27 88.32 96.00 96.53 96.00 96.10 54.72 48.01 54.73 48.45
Method Law Article Charge Prison Term
Acc Ma-P Ma-R Ma-F Acc Ma-P Ma-R Ma-F Acc Ma-P Ma-R Ma-F
CNN(LeCun et al., 1989) 76.14 35.48 38.55 35.39 74.91 74.00 78.12 73.97 27.38 18.48 17.51 17.44
BERT(Devlin et al., 2019) 82.62 45.89 47.91 45.83 80.50 80.34 81.09 78.36 36.80 29.83 27.50 27.03
Roberta(Zhong et al., 2018) 80.32 42.36 44.22 41.80 79.26 78.93 81.25 78.18 29.74 24.73 24.76 23.22
TopJudge(Zhong et al., 2018) 78.73 40.38 41.47 40.09 76.67 74.00 77.40 74.62 27.14 19.76 17.69 17.94
R-FormerDong and Niu (2021) 87.69 53.03 49.35 50.23 90.71 93.06 88.66 89.82 38.63 32.63 32.76 29.51
LADANXu et al. (2020) 79.44 48.43 44.13 46.18 79.64 48.43 44.13 46.18 33.69 26.40 22.94 24.55
Neurjudge(Yue et al., 2021) 71.38 52.86 53.52 52.62 71.85 69.37 71.09 68.66 26.80 26.81 26.85 25.97
EPMFeng et al. (2022) 84.19 47.21 43.79 44.39 83.49 80.36 83.29 81.87 36.91 30.65 31.61 30.20
CTM(Liu et al., 2022) 79.44 47.83 42.25 43.43 79.33 82.39 83.12 82.81 36.81 27.10 25.96 26.46
Llama3-chineseZhichen Zhang (2024) 3.01 2.21 1.09 1.36 27.72 37.12 27.72 28.28 15.66 19.18 15.66 15.56
PLJP(Bert)(Wu et al., 2023b) 94.18 74.65 76.23 74.84 94.18 90.25 88.67 89.05 43.52 33.37 35.67 31.98
D-LADAN(Xu et al., 2024) 90.65 44.95 48.45 44.49 88.95 29.41 27.64 28.48 46.18 24.80 23.78 22.73
RLJP(ours) 94.55 94.30 90.94 91.28 96.12 97.60 96.76 96.83 48.50 49.44 50.19 49.18
Table 2: Experiment results of LJP in CJO22 dataset. “Bold” indicates optimal results, and “underline” indicates sub-optimal results. The experimental results represent the average values obtained from five test rounds.

### v2

@@ 1 Introduction
The current methods of LJP mainly consist of two approaches: modeling legal knowledge and retrieving judicial precedents or legal knowledge. (1) Methods modeling legal knowledge typically use neural networks to model legal provisions and extract semantic associations between case facts and legal knowledgeXuWCPWZ20; RN31. While these models can efficiently extract relevant legal knowledge for LJP relying on semantic similarity, they overlook the intrinsic logic of cases. (2) Other methods leveraged semantic techniques to search for related judicial precedents or legal knowledgeRN33; RN35. Judicial precedents and legal knowledge as references to enhance the model’s understanding of the case circumstances, but they tend to study similar cases or knowledge instead of modeling the logical reasoning process. In summary, the above two types of methods rely on textual and semantic matching but ignore capturing the logic of judging legal cases, where logic in reasoning is crucial in legal judgment.
To address the aforementioned issues, researchers applied legal judgment logic to enhance reasoning in traditional deep learning models or large language models (LLMs) for LJP. Some researchers employ deep learning models to extract logical premises from complex cases Yue0JWZACYW21; RN38 or handle the legal judgment using rules defined by experts gan2021judgment. The integration of legal judgment logic has improved the reasoning capabilities of deep learning models. However, the ability of deep learning models is still limited by their model capacity and training data scale. As LLMs have remarkable abilities in text comprehension and logical reasoning, researchers combined judgment logic with LLMs for LJP. Researchersdeng2023syllogistic use LLMs to retrieve relevant legal provisions and extract four types of criminal elements from case facts for judgment based on the classical legal logic syllogism. Prompts constructed by keyword structure from legal experts’ determination guide LLMs to classify legal casesizzidien2024llm. In summary, while existing methods using legal logic effectively analyze cases through established rules, their logic rigidity limits adaptation to case-specific contexts in complex scenarios.
@@ 2 Related Work
LJP aims to predict judicial outcomes by employing traditional models or LLMs to analyze legal cases 10216994; RN30; zhang2024hd. Broadly, LJP can be divided into two main approaches: one based on semantic similarity liu2024semdr and the other guided by judgment logic cheng2024research.
@@ 2.1 LJP Assisted by Semantic Similarity
The semantic similarity-assisted method for LJP involves modeling domain knowledge and applying retrieval techniques zhang2023case.
The former mainly utilizes neural networks to model legal knowledge Yue0JWZACYW21; yao2024lawyer for feature extraction and to establish semantic associations among the knowledge. To improve the accuracy of crime prediction, the prevailing methods primarily focus on refining the attention mechanism 10216994; yu2023enhancing; he2025simulating. Additionally, some methods infuse domain-specific legal knowledge into LLMs wang2024legalreasoner; RN44, thereby enhancing the models’ capability to identify and leverage critical information. Regarding confusion issue in criminal charges, zhang2023contrastive (zhang2023contrastive) and gan-etal-2023-exploiting (gan-etal-2023-exploiting) utilized contrastive learning to capture fine-grained distinctions. Using retrieval techniques refers to utilizing search engines RN33; RN34; RN41 to obtain external knowledge from semantically similar judicial precedents. RN35 (RN35) introduced the retrieval technology into the SimCourt multi-agent framework. RN47 (RN47) constructed a hierarchical semantic ID for the retrieved candidate documents.
@@ 2.2 LJP Assisted by Case Judgment Logic
The legal judgment logic-enhanced method for LJP leverages models to summarize the knowledge embedded in case details chen2020legal; yang2022mve and extract the underlying reasoning logic.
Existing methods focus on using the entire factual description chai2020description; hong2023improving to generate judgment results, overlooking the actual judicial scenario where judges consider various criminal circumstances wu2022towards to determine the verdict and sentencing. Consequently, Yue0JWZACYW21 (Yue0JWZACYW21) proposed a legal judgment framework, NeurJudge, which can divide facts and the prediction results of intermediate subtasks into different scenarios. RN38 (RN38) further proposed NeurJudge+, which integrates the semantic labels of accusations and legal provisions into facts. gan2021judgment (gan2021judgment) injected legal knowledge as an additional logical rule into the neural network. deng2023syllogistic use LLMs to retrieve relevant legal provisions and extract four types of criminal elements from case facts for judgment based on the classical legal syllogism. Prompts constructed by keyword structure from legal experts’ determination guide LLMs to classify legal casesizzidien2024llm.
@@ 3.1 Task Definition
  • Judgment is the final decision made by a judge in a legal case based on the facts and the judgment rule. It typically consists of the law article, the charge, and the prison term. We represent the judgment of a case as j=(article,charge,prison_term)j=(article,charge,prison\_term), where (article,charge,prison_term)(article,charge,prison\_term) refers to the labels of provisions, accusation, and prison term, respectively.
@@
Problem 2 Legal Judgment Prediction. Given the case fact $f$, our task is to analyze $f$ based on rulerule, then predict judgment j=(article,charge,prison_term)j=(article,charge,prison\_term).
@@ 3.3 Rules Initialization
Each FOL judgment rule $Rule:A\to C$ comprises an antecedent $A$ and a consequent $C$. The content of $A$ is determined by the circumstance logic. And $C$ is constructed by one or two legal judgment labels, including articlearticle, (articlearticle, chargecharge), and (articlearticle, prison_termprison\_term).
@@
(3) Construct FOL judgment rules. The purpose of this step is to standardize the format of rules to facilitate subsequent optimization and reasoning processes. The judgment rules are constructed using the antecedent $A$ and the consequent $C$. $Rule:A\to C$ in this step. The antecedent $A$ comprises multiple FOL symbols and logical connectives ($\vee$, $\wedge$, and $\neg$). The consequent $C$ is the same as LJP labels, including legal provisions articlearticle, (legal provisions articlearticle, accusation chargecharge), and (legal provisions articlearticle, prison terms prison_termprison\_term).
@@ 3.4 Confusion-Aware Rule Optimize Engine
0:PrecedentsPrecedents: Precedent set;
0:targettarget: Target label which is the consequence of the rules;
0:RULERULE: Initial rule set ;
0:DefinedScoreDefinedScore: A predefined score threshold serves as the stopping criterion for terminating the iterative updating process;
0:NumNum: The required number of confusable cases.
[…]
@@
[…]
@@
9:  if Is_Evaluate($r$) then
10:   (scorescore, reasoning records) $\leftarrow history(r)$
@@
12:   (scorescore, reasoning records) $\leftarrow Evaluate(V_{target},r)$
@@
22:  $R^{\prime}\leftarrow CACL(R^{*},$ reasoning records ))
@@ 3.4.1 Construction of the Confusable Cases Validation Set
To find confusable cases for legal judgment targettarget, we first find the case set, whose legal judgment label is targettarget, denoted as $S_{positive}$, and the case set for other legal judgment labels, denoted as $S_{other}$. Then, to asses the similarity score between $S_{positive}$ and $S_{other}$, we convert each case fact set into fixed-dimensional embedding vectors using the BGE model: $E_{t}=\left\{E_{t}^{1},E_{t}^{2},...,E_{t}^{m}\right\}$ for $S_{positive}$ and $E_{o}=\left\{E_{o}^{1},E_{o}^{2},...,E_{o}^{n}\right\}$ for $S_{other}$, $E_{t}^{i},E_{o}^{i}\in R^{d}$. Then, we compute the cosine similarity between $E_{t}$ and $E_{o}$. The results are organized into a similarity matrix $A$, where $A(i,j)=\frac{E_{t}^{i}\cdot E_{o}^{i}}{\left\|E_{t}^{i}\right\|\left\|E_{o}^{i}\right\|}$ represents the similarity score of two cases $S_{positive}^{i}$ and $S_{other}^{j}$.
The higher the similarity score between cases in different judgment labels means that they contain similar crime circumstances, the more likely they are to be confusable for legal judgment. For this reason, we sort the similarity matrix $A$ in descending order based on the similarity scores of each row, and the first column is extracted to obtain a set of negative sample cases for judgment targettarget: $S_{negative}\in S_{other}$. Finally, the confusable case set $V_{target}$ can be constructed: $V_{target}=S_{negative}\cup S_{positive}$.
@@ 3.4.2 Optimization Tree Splitting
[…]
@@ 4.1 Settings
Datasets. Following previous LJP works (ZhongGTX0S18; XuWCPWZ20; Yue0JWZACYW21; WuZL0LZS0K23), we conduct experiments on datasets CAIL2018  [^2] https://github.com/china-ai-law-challenge and CJO22(WuZL0LZS0K23). For the CAIL2018 dataset, we randomly divide it into training, validation, and test sets with a ratio of 8:1:1WuZL0LZS0K23. For the CJO22 dataset, following WuZL0LZS0K23, we use it solely as an additional test set. The table in Appendix F presents the details of two datasets.
Evaluation Metrics. We use Accuracy (Acc), Macro-Precision (Ma-P), Macro-Recall (Ma-R), and Macro-F1 (Ma-F) as the evaluation metrics following previous LJP works (ZhongGTX0S18; XuWCPWZ20; Yue0JWZACYW21; WuZL0LZS0K23).
Baselines. We compare with: (1) CNN(LeCunBDHHHJ89) use different kernel convolutional operations to extract text features for classification; (2) BERT(DevlinCLT19) can be easily fine-tuned on downstream tasks such as LJP; (3) TopJudge(ZhongGTX0S18) employs multi-task learning and captures the dependencies among the three sub-tasks in LJP; (4) R-FormerDongN21 formalizes LJP as a node classification problem over a global consistency graph; (5) LADANXuWCPWZ20 uses graph distillation to extract discriminative features of the fact; (6) NeurJudge(Yue0JWZACYW21) splits the fact description into different parts for making predictions; (7) EPMDBLP:conf/ijcai/FengL022 locates event-related information essential for judgment; (8) CTM(LiuDLP022) uses case triple modeling from contrastive case relations; (9) PLJP(WuZL0LZS0K23) uses the collaboration between LLMs and domain-specific models for LJP; (10) Llama3(WuZL0LZS0K23), which is Meta’s advanced open-source language model; (11) D-LADAN(RN31) solves LJP confusion via graph and memory mechanisms.
Implementation Details. We directly adopted Bert-base-chinese [^3] https://huggingface.co/google-bert/bert-base-chinese for candidate labels, Gpt-4o for rules optimization, and Llama3-chineseLlama3-Chinese for other steps because the language of the cases in the data set is Chinese. For the length limit, we use three precedents in Section 3.3. For the parameter settings of the baseline, we strictly follow the original paper. More details are in Appendix G.
Method Law Article Charge Term-of-Penalty
Acc Ma-F Acc Ma-F Acc Ma-F
Navie Model BERT 82.77 35.82 89.1 89.63 40.00 33.58
Llama3-Chinese 2.62 1.11 22.63 10.04 12.00 10.53
Semantic-based Neurjudge 76.91 53.56 82.13 82.36 33.53 36.53
PLJP_Bert 87.07 56.63 94.99 91.33 48.72 35.43
D-LADAN 91.05 58.38 90.82 62.08 38.31 25.00
KnowJudge 89.00 85.00
Logic-based EPM 85.80 47.32 91.20 90.46 40.25 37.34
CTM 84.72 45.10 90.28 86.30 39.56 37.84
Ours(RLJP) 88.12 54.42 96.00 96.10 54.72 48.45
Table 2: Experiment results of LJP in CJO22 dataset. Bold and underlined denote the same notation as Tab. 1.
Law Article Charge Term-of-Penalty
Method Acc Ma-F Acc Ma-F Acc Ma-F
BERT 82.62 45.83 80.5 78.36 36.8 27.03
Navie model Llama3-Chinese 3.01 1.36 27.72 28.28 15.66 15.56
Neurjudge 71.38 52.62 71.85 68.66 26.8 25.97
PLJP_Bert 94.18 74.84 94.18 89.05 43.52 31.98
Semantic-based D-LADAN 90.65 44.49 88.95 28.48 46.18 22.73
EPM 84.19 44.39 83.49 81.87 36.91 30.2
Logic-based CTM 79.44 43.43 79.33 82.81 36.81 26.46
Ours(RLJP) 94.55 91.28 96.12 96.83 48.50 49.18